csc^2(x)−cot^2(x)=
1
Expert's answer
2020-11-16T13:47:32-0500

"\csc^2 x-\cot^2 x=\frac{1}{\sin^2 x}-\frac{\cos^2 x}{\sin^2 x}=\frac{1-\cos^2 x}{\sin^2 x}=\frac{\sin^2 x}{\sin^2 x}=1"